OpenStudy (anonymous):

A line has a vertical change per unit of horizontal change of negative two-fifths and passes through point one comma one. Which is another point on the line?

OpenStudy (anonymous):

\[y-1=-\frac{ 2 }{ 5 }(x-1)\]

OpenStudy (anonymous):

use this equation to generate as many points as you like: \[y=-\frac{ 2 }{ 5 }x+\frac{ 7 }{ 5 }\]

OpenStudy (anonymous):

if x = 0 then y = (7/5)